OpenStudy (anonymous):

Explain why the quantum number set (2, 1, -2, -½) is not possible for an electron in an atom.?

OpenStudy (chmvijay):

Bcoz the values n=2, l=1 and ml= +1 .O -1 and S =-1/2 since in this case ml= -2 is not possible for l=1 case